when a single card is drawn from an ordinary 52-card deck, find the odds in favor of getting a red 6 or a black 2
Answer by edjones(8007)

2/52 + 2/52 = 4/52 = 1/13 probability that a red 6 or a black 2 will be drawn.
12/13 probability that a red 6 or a black 2 will not be drawn.
1 to 12 Odds in favor of getting a red 6 or a black 2. (the numerators of the two fractions)
.
Ed